# One-Prompt Strikes Back: Sparse Mixture of Experts for Prompt-based Continual Learning

### v1

@@ Appendix B Training Algorithm of SMoPE
Following prior work on SMoE training [48, 25, 4], we first adopt dense training (K=1K=-1) for several initial epochs on the first task. This provides a good initialization of the prompt parameters before enabling sparse expert selection.
For each task $t$, the model is optimized on $\mathcal{D}_{t}$ using the objective in Equation (18). We then refine the classifier parameters $\phi$ using the task-adaptive prediction (TAP) objective, following [45, 18, 23], which mitigates classifier bias by modeling the Gaussian distributions of previously seen classes. Specifically, for each class $c\in\mathcal{Y}^{i}$ from tasks $i=1,\dots,t-1$, we maintain a Gaussian distribution $\mathcal{G}_{c}=\mathcal{N}\left(\bm{\mu}_{c},\bm{\Sigma}_{c}\right)$, approximated from the prompted representations $\mathcal{D}_{c}^{z}=\left\{{\bm{z}}=f_{\theta}({\bm{x}};\mathbf{P},K)\ |\ ({\bm{x}},c)\in\mathcal{D}_{i}\right\}$. The distribution parameters are estimated as:
@@ Appendix C Additional Experimental Details
Evaluation Metrics. We report two standard metrics, Final Average Accuracy (FAA) and Cumulative Average Accuracy (CAA), as they inherently capture both plasticity and forgetting [38]. FAA measures the overall performance by computing the average accuracy across all $T$ tasks after the completion of continual learning. CAA extends this by averaging the performance across all intermediate stages. Formally, let $S_{i,t}$ denote the accuracy on the $i$-th task after learning the $t$-th task, and define the average accuracy after $t$ tasks as $A_{t}=\frac{1}{t}\sum_{i=1}^{t}S_{i,t}$. Then, after all $T$ tasks are learned, FAA and CAA are given by $\mathrm{FAA}=A_{T}$, and $\mathrm{CAA}=\frac{1}{T}\sum_{t=1}^{T}A_{t}$.
Data Augmentation. Input images are resized to $224\times 224$ and augmented following the protocol of [38], including random horizontal flipping and standard normalization.
Implementation Details. We use a pre-trained ViT-B/16 model as the backbone. Training is conducted using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. Following the setup in VQ-Prompt [18], 20% of the training data is reserved for validation, and a grid search is used to tune hyperparameters. We evaluate L2P++, Deep L2P++, DualPrompt, CODA-Prompt, and VQ-Prompt using the official implementation provided by VQ-Prompt. For OVOR, we similarly employ the original codebase. Recent work by [12] has identified an implementation issue in the official HiDe-Prompt codebase, specifically in its prompt retrieval mechanism. This issue introduces information leakage, resulting in overestimated performance, particularly when inference is performed with a batch size of 1, while the training batch size remains consistent with the original setup. To address this, we adopt the corrected prompt retrieval code provided by [12] for both HiDe-Prompt and NoRGa, while preserving all other components from the original implementations.
@@ D.1 Performance Across Varying Task Lengths
Method 5-task 10-task 20-task
Joint-Train 82.06 82.06 82.06
L2P++ $70.83\pm 0.58$ $78.34\pm 0.47$ $69.29\pm 0.73$ $78.30\pm 0.69$ $65.89\pm 1.30$ $77.15\pm 0.65$
Deep L2P++ $73.93\pm 0.37$ $80.14\pm 0.54$ $71.66\pm 0.64$ $79.63\pm 0.90$ $68.42\pm 1.20$ $78.68\pm 1.03$
DualPrompt $73.05\pm 0.50$ $79.47\pm 0.40$ $71.32\pm 0.62$ $78.94\pm 0.72$ $67.89\pm 1.39$ $77.42\pm 0.80$
CODA-Prompt $76.51\pm 0.38$ $82.04\pm 0.54$ $75.45\pm 0.56$ $81.59\pm 0.82$ $72.37\pm 1.19$ $79.88\pm 1.06$
OVOR $75.81\pm 0.16$ $79.31\pm 0.42$ $75.25\pm 0.21$ $79.78\pm 0.65$ $2.45\pm 0.42 77$$7.60\pm 1.15$
HiDe-Prompt $75.00\pm 0.05$ $79.68\pm 0.43$ $74.25\pm 0.19$ $79.64\pm 0.53$ $73.71\pm 0.37$ $79.14\pm 0.72$
NoRGa $75.17\pm 0.19$ $79.64\pm 0.62$ $74.39\pm 0.08$ $79.68\pm 0.41$ $73.89\pm 0.33$ $79.22\pm 0.84$
VQ-Prompt $79.23\pm 0.29$ $82.96\pm 0.50$ $78.71\pm 0.22$ $83.24\pm 0.68$ $\textbf{78.10}\pm 0.22$ $82.70\pm 1.16$
SMoPE $\textbf{80.09}\pm 0.42$ $\textbf{84.47}\pm 0.95$ $\textbf{79.32}\pm 0.42$ $\textbf{84.39}\pm 0.77$ $77.81\pm 0.36$ $\textbf{83.38}\pm 0.98$
To assess the scalability and generalizability of SMoPE, we follow prior work [15, 18] by partitioning ImageNet-R into 5, 10, and 20 sequential tasks. The results are shown in Table 6. Across all configurations, SMoPE consistently outperforms task-specific methods such as HiDe-Prompt and NoRGa, despite using a single shared prompt. Its performance remains comparable to VQ-Prompt, further validating its robustness under varying task lengths. These findings suggest that SMoPE can generalize effectively as the number of tasks increases. However, relying on a single prompt for an indefinite number of tasks may eventually lead to knowledge interference as the task sequence grows. To address this limitation, future work could explore strategies for scaling the number of prompt experts to better support longer task sequences.
@@ D.3 Detailed Analysis of Adaptive Noise
Previous work has typically encouraged expert exploration by injecting random noise into expert scores, either with a fixed magnitude [32] or by sampling from predefined distributions (e.g., Uniform) [37, 11]. However, in pre-trained models, the range of expert scores can vary considerably across attention heads (see Figure 4), making it difficult to select an appropriate fixed noise level. In contrast, our method requires tuning only a single noise parameter, $\epsilon$, which is constrained to the interval $[0,1]$. This approach eliminates the need for head-specific noise scaling and significantly reduces the complexity of hyperparameter tuning in large-scale models.
@@ D.5 Detailed Analysis of Prototype Loss
This perspective aligns with findings in recent literature. For example, [6] showed that routers can learn cluster-center features, facilitating the decomposition of complex tasks into simpler sub-problems addressable by individual experts. Furthermore, the expert selection process can be interpreted as a classification problem, where the prefix keys act as a linear classification head. In this context, recent studies have proposed that the weight vector associated with each class in a classifier can be interpreted as a class prototype [39, 33, 52], further supporting our hypothesis.
@@ D.6 Sample Efficiency Evaluation
As demonstrated in Appendix A, applying the prompt-attention score aggregation technique yields MoE models within attention heads that maintain the same estimation rate and sample efficiency as those achieved by the standard prefix tuning formulation. To empirically support this theoretical result, we follow [23] and evaluate sample efficiency by tracking the validation loss on ImageNet-R throughout the training process for each new task. The results, presented in Figure 9, show that both models exhibit comparable convergence rates. This suggests that incorporating prompt-attention score aggregation does not compromise the model’s ability to efficiently adapt to new tasks.

### v2

@@ Appendix B Training Algorithm of SMoPE
Following prior work on SMoE training (Wu et al., 2022; Lin et al., 2024; Cai et al., 2025), we first adopt dense training (K=1K=-1) for several initial epochs on the first task. This provides a good initialization of the prompt parameters before enabling sparse expert selection.
For each task $t$, the model is optimized on $\mathcal{D}_{t}$ using the objective in Equation (18). We then refine the classifier parameters $\phi$ using the task-adaptive prediction (TAP) objective, following (Wang et al., 2023; Jiao et al., 2024; Le et al., 2024), which mitigates classifier bias by modeling the Gaussian distributions of previously seen classes. Specifically, for each class $c\in\mathcal{Y}^{i}$ from tasks $i=1,\dots,t-1$, we maintain a Gaussian distribution $\mathcal{G}_{c}=\mathcal{N}\left(\bm{\mu}_{c},\bm{\Sigma}_{c}\right)$, approximated from the prompted representations $\mathcal{D}_{c}^{z}=\left\{{\bm{z}}=f_{\theta}({\bm{x}};\mathbf{P},K)\ |\ ({\bm{x}},c)\in\mathcal{D}_{i}\right\}$. The distribution parameters are estimated as:
@@ Appendix C Additional Experimental Details
Evaluation Metrics. We report two standard metrics, Final Average Accuracy (FAA) and Cumulative Average Accuracy (CAA), as they inherently capture both plasticity and forgetting (Smith et al., 2023). FAA measures the overall performance by computing the average accuracy across all $T$ tasks after the completion of continual learning. CAA extends this by averaging the performance across all intermediate stages. Formally, let $S_{i,t}$ denote the accuracy on the $i$-th task after learning the $t$-th task, and define the average accuracy after $t$ tasks as $A_{t}=\frac{1}{t}\sum_{i=1}^{t}S_{i,t}$. Then, after all $T$ tasks are learned, FAA and CAA are given by $\mathrm{FAA}=A_{T}$, and $\mathrm{CAA}=\frac{1}{T}\sum_{t=1}^{T}A_{t}$.
Data Augmentation. Input images are resized to $224\times 224$ and augmented following the protocol of Smith et al. (2023), including random horizontal flipping and standard normalization.
Implementation Details. We use a pre-trained ViT-B/16 model as the backbone. Training is conducted using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. Following the setup in VQ-Prompt (Jiao et al., 2024), 20% of the training data is reserved for validation, and a grid search is used to tune hyperparameters. We evaluate L2P++, Deep L2P++, DualPrompt, CODA-Prompt, and VQ-Prompt using the official implementation provided by VQ-Prompt. For OVOR, we similarly employ the original codebase. Recent work by Feng et al. (2025) has identified an implementation issue in the official HiDe-Prompt codebase, specifically in its prompt retrieval mechanism. This issue introduces information leakage, resulting in overestimated performance, particularly when inference is performed with a batch size of 1, while the training batch size remains consistent with the original setup. To address this, we adopt the corrected prompt retrieval code provided by Feng et al. (2025) for both HiDe-Prompt and NoRGa, while preserving all other components from the original implementations.
@@ D.1 Performance Across Varying Task Lengths
Method 5-task 10-task 20-task 50-task
FAA ($\uparrow$) CAA ($\uparrow$) FAA ($\uparrow$) CAA ($\uparrow$) FAA ($\uparrow$) CAA ($\uparrow$) FAA ($\uparrow$) CAA ($\uparrow$)
Joint-Train 82.06 82.06 82.06 82.06
L2P++ $70.83\pm 0.58$ $78.34\pm 0.47$ $69.29\pm 0.73$ $78.30\pm 0.69$ $65.89\pm 1.30$ $77.15\pm 0.65$ $55.65\pm 0.89$ $64.88\pm 0.86$
Deep L2P++ $73.93\pm 0.37$ $80.14\pm 0.54$ $71.66\pm 0.64$ $79.63\pm 0.90$ $68.42\pm 1.20$ $78.68\pm 1.03$ $58.48\pm 0.54$ $67.04\pm 0.80$
DualPrompt $73.05\pm 0.50$ $79.47\pm 0.40$ $71.32\pm 0.62$ $78.94\pm 0.72$ $67.89\pm 1.39$ $77.42\pm 0.80$ $58.31\pm 1.18$ $65.51\pm 0.95$
CODA-Prompt $76.51\pm 0.38$ $82.04\pm 0.54$ $75.45\pm 0.56$ $81.59\pm 0.82$ $72.37\pm 1.19$ $79.88\pm 1.06$ $66.96\pm 0.71$ $75.19\pm 0.72$
OVOR $75.81\pm 0.16$ $79.31\pm 0.42$ $75.25\pm 0.21$ $79.78\pm 0.65$ $2.45\pm 0.42 77$$7.60\pm 1.15$ $65.84\pm 0.59$ $74.84\pm 0.69$
HiDe-Prompt $75.00\pm 0.05$ $79.68\pm 0.43$ $74.25\pm 0.19$ $79.64\pm 0.53$ $73.71\pm 0.37$ $79.14\pm 0.72$ $70.82\pm 0.54$ $76.81\pm 0.83$
NoRGa $75.17\pm 0.19$ $79.64\pm 0.62$ $74.39\pm 0.08$ $79.68\pm 0.41$ $73.89\pm 0.33$ $79.22\pm 0.84$ $71.02\pm 0.63$ $76.99\pm 0.58$
VQ-Prompt $79.23\pm 0.29$ $82.96\pm 0.50$ $78.71\pm 0.22$ $83.24\pm 0.68$ $\textbf{78.10}\pm 0.22$ $82.70\pm 1.16$ $75.31\pm 0.61$ $81.52\pm 0.81$
SMoPE $\textbf{80.09}\pm 0.42$ $\textbf{84.47}\pm 0.95$ $\textbf{79.32}\pm 0.42$ $\textbf{84.39}\pm 0.77$ $77.81\pm 0.36$ $\textbf{83.38}\pm 0.98$ $\textbf{75.54}\pm 0.49$ $\textbf{81.94}\pm 0.93$
To assess the scalability and generalizability of SMoPE, we follow prior work (Huang et al., 2024; Jiao et al., 2024) by partitioning ImageNet-R into 5, 10, and 20 sequential tasks, ensuring a fair comparison with existing methods. Additionally, we introduce a 50-task split to more rigorously evaluate SMoPE’s capacity to handle longer task sequences. The results are shown in Table 6. Across all configurations, SMoPE consistently outperforms task-specific methods such as HiDe-Prompt and NoRGa, despite using a single shared prompt. Its performance remains comparable to VQ-Prompt, further validating its robustness under varying task lengths. These findings suggest that SMoPE can generalize effectively as the number of tasks increases. However, relying on a single prompt for an indefinite number of tasks may eventually lead to knowledge interference as the task sequence grows. To address this limitation, future work could explore strategies for scaling the number of prompt experts to better support longer task sequences.
@@ D.3 Detailed Analysis of Adaptive Noise
Previous work has typically encouraged expert exploration by injecting random noise into expert scores, either with a fixed magnitude (Nguyen et al., 2024) or by sampling from predefined distributions (e.g., Uniform) (Shazeer et al., 2017; Fedus et al., 2022). However, in pre-trained models, the range of expert scores can vary considerably across attention heads (see Figure 4), making it difficult to select an appropriate fixed noise level. In contrast, our method requires tuning only a single noise parameter, $\epsilon$, which is constrained to the interval $[0,1]$. This approach eliminates the need for head-specific noise scaling and significantly reduces the complexity of hyperparameter tuning in large-scale models.
@@ D.5 Detailed Analysis of Prototype Loss
This perspective aligns with findings in recent literature. For example, Chen et al. (2022) showed that routers can learn cluster-center features, facilitating the decomposition of complex tasks into simpler sub-problems addressable by individual experts. Furthermore, the expert selection process can be interpreted as a classification problem, where the prefix keys act as a linear classification head. In this context, recent studies have proposed that the weight vector associated with each class in a classifier can be interpreted as a class prototype (Snell et al., 2017; Qi et al., 2018; Zhao et al., 2024), further supporting our hypothesis.
@@ D.6 Sample Efficiency Evaluation
As demonstrated in Appendix A, applying the prompt-attention score aggregation technique yields MoE models within attention heads that maintain the same estimation rate and sample efficiency as those achieved by the standard prefix tuning formulation. To empirically support this theoretical result, we follow Le et al. (2024) and evaluate sample efficiency by tracking the validation loss on ImageNet-R throughout the training process for each new task. The results, presented in Figure 9, show that both models exhibit comparable convergence rates. This suggests that incorporating prompt-attention score aggregation does not compromise the model’s ability to efficiently adapt to new tasks.
D.7 Analysis of Prompt Expert Activation Frequencies
Figure 10: Evolution of Prompt Expert Activation Frequencies Across Sequential Tasks. Results on CUB-200 with a prompt length of $N_{p}=25$ and $K=5$. We present the results for a representative attention head and visualize the frequency with which prompt experts are activated across sequential tasks.
To provide direct evidence for our central claim that sparse activation mitigates interference, this section analyzes how expert activation patterns evolve as the model encounters a sequence of tasks. As illustrated in Figure 10, which visualizes activation frequencies on the CUB-200 dataset, the model initially activates a sparse and concentrated subset of prompt experts. As training progresses across subsequent tasks, the set of activated experts becomes increasingly larger and more diverse.
This dynamic and adaptive activation pattern is fundamental to how our method mitigates interference. In contrast to dense, single-prompt baselines (e.g., OVOR), where a single set of parameters must be continually updated for each new task—risking the overwriting of previously acquired knowledge—our sparse approach distributes the learning load across a broader pool of specialized experts. New tasks can activate and train new or underutilized experts, leaving the parameters of experts essential for earlier tasks largely unaffected. This selective activation is crucial for preserving acquired knowledge while maintaining plasticity for new learning.
Furthermore, this diversification does not preclude effective knowledge reuse. The activation mechanism in the SMoPE framework is input-driven rather than task-ID-driven, as discussed in Section 3.2. Consequently, an expert can be reactivated for subsequent tasks if its specialized function is relevant to the current input. This mechanism facilitates positive knowledge transfer while preventing the destructive interference.
D.8 Accuracy Curves Across Incremental Steps
To provide a more comprehensive understanding of the method’s behavior, we present accuracy curves illustrating how FAA performance evolves across sequential tasks on CUB-200 under the 10-task setting. The results are shown in Figure 11. As observed, OVOR performance deteriorates rapidly as training progresses. This decline can be attributed to knowledge interference arising from updating all prompt parameters simultaneously. In contrast, SMoPE maintains consistently strong performance across all tasks. Notably, its accuracy surpasses that of methods employing task-specific prompts, such as HiDe-Prompt. These results demonstrate the effectiveness and robustness of the SMoPE architecture, even when using shared prompt parameters.
Figure 11: Accuracy Curves Across Incremental Steps. Performance of different methods on CUB-200 under the 10-task incremental learning setting.
Appendix E Use of Large Language Models
Large language models were employed solely for editorial purposes, including grammar correction and spelling refinement. They were not used for content generation, data analysis, or the design of experiments.